# EPO: Entropy-regularized Policy Optimization for LLM Agents Reinforcement Learning

### v1

@@ A.2 The EPO Performance Bound
Expanding the EPO objectives:
$V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})$ $\geq V_{\lambda,\beta}^{\pi^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})$ (8)
$=\left(V^{\pi^{*}}(s_{0})+\lambda\mathcal{H}(\pi^{*}|s_{0})-\lambda\beta_{k}\mathbb{E}_{\pi^{*}}[L^{\text{smooth}}|s_{0}]\right)$
$\quad-\left(V^{\pi_{\theta}}(s_{0})+\lambda\mathcal{H}(\pi_{\theta}|s_{0})-\lambda\beta_{k}\mathbb{E}_{\pi_{\theta}}[L^{\text{smooth}}|s_{0}]\right)$ (9)
Rearranging:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})$ $\leq V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})$
$\quad+\lambda\left(\mathcal{H}(\pi_{\theta}|s_{0})-\mathcal{H}(\pi^{*}|s_{0})\right)$
$\quad-\lambda\beta_{k}\left(\mathbb{E}_{\pi_{\theta}}[L^{\text{smooth}}|s_{0}]-\mathbb{E}_{\pi^{*}}[L^{\text{smooth}}|s_{0}]\right)$ (10)
We use two fundamental bounds on policy entropy. First, the entropy of an optimal policy $\pi^{*}$ is bounded by the maximum possible entropy over its support set, the optimal trajectories $\mathcal{A}_{H}^{*}(s_{0})$. Second, the entropy of any policy $\pi_{\theta}$ is bounded by the sum of the maximum possible single-step entropies over the horizon $H$.
$\mathcal{H}_{\max}$ $=-\sum_{\tau\in\mathcal{A}_{H}^{*}(s_{0})}\frac{1}{|\mathcal{A}_{H}^{*}(s_{0})|}\log\frac{1}{|\mathcal{A}_{H}^{*}(s_{0})|}$
$=-|\mathcal{A}_{H}^{*}(s_{0})|\left(\frac{1}{|\mathcal{A}_{H}^{*}(s_{0})|}(-\log|\mathcal{A}_{H}^{*}(s_{0})|)\right)$
$=\log|\mathcal{A}_{H}^{*}(s_{0})|$
$\mathcal{H}(\pi_{\theta}|s_{0})$ $=\mathcal{H}(a_{0},a_{1},\dots,a_{H-1}|s_{0})$
$=\sum_{t=0}^{H-1}\mathcal{H}(a_{t}|s_{t})$
We then obtained bounds showing that the entropy of the optimal policy concentrates on optimal actions:
$\mathcal{H}(\pi^{*}|s_{0})\leq\log|\mathcal{A}_{H}^{*}(s_{0})|$
and established a maximum entropy bound:
$\mathcal{H}(\pi_{\theta}|s_{0})\leq H\log|\mathcal{A}|.$
@@
$\mathcal{H}(\pi_{\theta}|s_{0})-\mathcal{H}(\pi^{*}|s_{0})\leq H\log|\mathcal{A}|-\log|\mathcal{A}_{H}^{*}(s_{0})|=H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}$
This gives us:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}-\lambda\beta_{k}\Phi^{\pi_{\theta},\pi^{*}}(s_{0})$ (11)
We adapt the proof technique from Lemma 1. The key modification is that our Q-function now includes the smoothing term. Given the deterministic transitions in our LLM setting (where $s_{t+1}$ is fully determined by $(s_{t},a_{t})$), we have:
$Q_{t,\lambda,\beta}^{\pi_{\theta}}(s,a)=r(s,a)+V_{t+1,\lambda,\beta}^{\pi_{\theta}}(s_{t+1})$
where the value function incorporates smoothing penalties. Following the derivation in the reference (equations leading to their Lemma 4), we express the performance gap as:
$V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})\leq\sum_{t=0}^{H-1}\mathbb{E}_{s\sim\mathbb{P}_{t}^{\pi_{\lambda,\beta}^{*}}(\cdot|s_{0})}\left[\lambda D_{KL}(\pi_{\theta}(\cdot|s)||\bar{\pi}_{\theta,\beta}(\cdot|s))\right]$
where $\bar{\pi}_{\theta,\beta}(a|s)\propto\exp(Q_{t,\lambda,\beta}^{\pi_{\theta}}(s,a)/\lambda)$ is the soft-optimal policy.
The KL divergence can be bounded using the softmax parameterization:
$D_{KL}(\pi_{\theta}(\cdot|s)||\bar{\pi}_{\theta,\beta}(\cdot|s))\leq\frac{1}{2\lambda^{2}}\left\|Q_{t,\lambda,\beta}^{\pi_{\theta}}(s,\cdot)-\lambda\theta_{s,\cdot}-c(s)\mathbf{1}\right\|_{\infty}^{2}$
where $c(s)$ is a normalizing constant and $\mathbf{1}$ is the all-ones vector.
The gradient of the EPO objective is:
$\nabla V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D})=\sum_{t=0}^{H-1}\sum_{s}\mathbb{P}_{t}^{\pi_{\theta}}(s)\frac{\partial\pi_{\theta}(\cdot|s)}{\partial\theta_{s,\cdot}}\left(Q_{t,\lambda,\beta}^{\pi_{\theta}}(s,\cdot)-\lambda\theta_{s,\cdot}\right)$
Let us define the soft advantage term $\bar{A}_{t,\lambda,\beta}^{\pi_{\theta}}(s)$ as:
$\bar{A}_{t,\lambda,\beta}^{\pi_{\theta}}(s):=Q_{t,\lambda,\beta}^{\pi_{\theta}}(s,\cdot)/\lambda-\theta_{s,\cdot}-c(s)\mathbf{1}$ (12)
where $c(s)$ is a state-dependent normalizing constant. Adapting the bounding technique from the reference, which combines the Cauchy-Schwarz inequality with properties of the softmax Jacobian, yields the following lower bound on the squared gradient norm:
$\|\nabla V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D})\|^{2}\geq\lambda^{2}C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})\sum_{t=0}^{H-1}\mathbb{E}_{s\sim\mathbb{P}_{t}^{\pi_{\lambda,\beta}^{*}}(\cdot|s_{0})}\left\|\bar{A}_{t,\lambda,\beta}^{\pi_{\theta}}(s)\right\|_{\infty}^{2}$ (13)
This inequality is crucial as it connects the global gradient norm to the sum of local, per-state soft advantages.
where:
$C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})=C_{d}^{2}\min_{t,s\in\mathcal{S}(s_{0})}\mathbb{P}_{t}^{\pi_{\theta}}(s)(\min_{s,a}\pi_{\theta}(a|s))^{2}\min_{t,s\in\mathcal{S}(s_{0})}\frac{\mathbb{P}_{t}^{\pi_{\theta}}(s)}{\mathbb{P}_{t}^{\pi_{\lambda,\beta}^{*}}(s|s_{0})}$
with $C_{d}=(|\mathcal{A}|^{H})^{-1/2}$ and $\mathcal{S}(s_{0})$ being the set of all states reachable from $s_{0}$.
@@
$V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}(s_{0})-V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})\leq\frac{|\mathcal{D}|^{2}}{2\lambda C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}\|\nabla V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D})\|^{2}$ (14)
Substituting Equation 14 into Equation 11, and using the assumption $\|\nabla V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D})\|\leq\epsilon$:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{|\mathcal{D}|^{2}}{C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}-\lambda\beta_{k}\Phi^{\pi_{\theta},\pi^{*}}(s_{0})$
The bias reduction term $\lambda\beta_{k}\Phi^{\pi_{\theta},\pi^{*}}(s_{0})$ effectively counteracts standard entropy bias when the optimal policy exhibits stable, low-variance entropy and the current policy experiences entropy violations. This correction becomes increasingly effective through the dynamic coefficient $\beta_{k}$, which strengthens as training progresses to enable smooth transition from exploration-focused early training to stability-focused convergence. Consequently, EPO achieves strictly better bounds than standard maximum entropy RL when $\beta_{k}\Phi^{\pi_{\theta},\pi^{*}}(s_{0})>0$, directly addressing the large bias issue that becomes prohibitive in LLM settings.
A.3 Core Proof Steps for Lemmas
Lemma 1 (Performance Difference). For any $h\in\{0,1,\dots,H-1\}$ and state $s\in\mathcal{S}$, the performance difference between any two policies $\pi$ and $\pi^{\prime}$ is: V_h^π(s)-V_h^π’(s)=E_π[_t=h^H-1 A_t^π’(s_t,a_t)|s_h=s]
Proof.
$V_{h}^{\pi}(s)-V_{h}^{\pi^{\prime}}(s)$ $=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H-1}r(s_{t},a_{t})\bigg|s_{h}=s\right]-V_{h}^{\pi^{\prime}}(s)$ (15)
$=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H-1}r(s_{t},a_{t})+\sum_{t=h}^{H-2}V_{t+1}^{\pi^{\prime}}(s_{t+1})-\sum_{t=h}^{H-2}V_{t+1}^{\pi^{\prime}}(s_{t+1})\bigg|s_{h}=s\right]-V_{h}^{\pi^{\prime}}(s)$ (16)
$=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H-1}Q_{t}^{\pi^{\prime}}(s_{t},a_{t})-\sum_{t=h}^{H-1}V_{t}^{\pi^{\prime}}(s_{t})\bigg|s_{h}=s\right]$ (17)
$=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H-1}A_{t}^{\pi^{\prime}}(s_{t},a_{t})\bigg|s_{h}=s\right]$ (18)
The key insight is telescoping the value functions and using $Q_{t}^{\pi^{\prime}}(s_{t},a_{t})=r(s_{t},a_{t})+V_{t+1}^{\pi^{\prime}}(s_{t+1})$.
Lemma 2 (Entropy Bias). It holds that: V^π^*(s_0)- V^π_θ(s_0) V_λ^π_λ^*(s_0)- V_λ^π_θ(s_0)+λHlog|A||AH*(s0)|1H where $\pi_{\lambda}^{*}=\arg\max_{\pi}V_{\lambda}^{\pi}(\mathcal{D})$ and $\pi^{*}\in\arg\max_{\pi}V^{\pi}(\mathcal{D})$.
By optimality of $\pi_{\lambda}^{*}$ for the entropy-regularized objective:
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})$ $\geq V_{\lambda}^{\pi^{*}}(s_{0})$ (19)
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})$ $\geq V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})+\lambda(\mathcal{H}(\pi^{*}|s_{0})-\mathcal{H}(\pi_{\theta}|s_{0}))$ (20)
The key bounds on entropy:
$\mathcal{H}(\pi^{*}|s_{0})$ $\leq\log|\mathcal{A}_{H}^{*}(s_{0})|\quad\text{(optimal policy is concentrated)}$ (21)
$\mathcal{H}(\pi_{\theta}|s_{0})$ $\leq H\log|\mathcal{A}|\quad\text{(maximum entropy bound)}$ (22)
Therefore:
$\mathcal{H}(\pi^{*}|s_{0})-\mathcal{H}(\pi_{\theta}|s_{0})\geq\log|\mathcal{A}_{H}^{*}(s_{0})|-H\log|\mathcal{A}|=-H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}$
Lemma 3 (Performance Bound under Gradient Norm). Assume the policy is softmax. If $\|\nabla V_{\lambda}^{\pi_{\theta}}(\mathcal{D})\|\leq\epsilon$: V_λ^π_λ^*(s_0)- V_λ^π_θ(s_0) 1|D|2Cλπθ(s0)ϵ^2
First, express the performance gap using KL divergence:
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})$ $\leq\sum_{t=0}^{H-1}\mathbb{E}_{s\sim\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(\cdot|s_{0})}\left[\lambda D_{KL}(\pi_{\theta}(\cdot|s)||\bar{\pi}_{\theta}(\cdot|s))\right]$ (23)
where $\bar{\pi}_{\theta}(a|s)\propto\exp(Q_{t,\lambda}^{\pi_{\theta}}(s,a)/\lambda)$.
For softmax policies, the KL divergence bound:
$D_{KL}(\pi_{\theta}(\cdot|s)||\bar{\pi}_{\theta}(\cdot|s))\leq\frac{1}{2\lambda^{2}}\left\|Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)/\lambda-\theta_{s,\cdot}-c(s)\mathbf{1}\right\|_{\infty}^{2}$
The gradient norm lower bound (using Cauchy-Schwarz):
$\|\nabla V_{\lambda}^{\pi_{\theta}}(\mathcal{D})\|^{2}$ $\geq C_{d}^{2}\sum_{s\in\mathcal{S}(s_{0})}\sum_{t=0}^{H-1}(\mathbb{P}_{t}^{\pi_{\theta}}(s))^{2}(\min_{a}\pi_{\theta}(a|s))^{2}$ (24)
$\quad\times\left\|Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)/\lambda-\theta_{s,\cdot}-c(s)\mathbf{1}\right\|_{\infty}^{2}$ (25)
Combining these with the constant:
$C_{\lambda}^{\pi_{\theta}}(s_{0})=C_{d}^{2}\min_{t,s}\mathbb{P}_{t}^{\pi_{\theta}}(s)(\min_{a}\pi_{\theta}(a|s))^{2}\min_{t,s}\frac{\mathbb{P}_{t}^{\pi_{\theta}}(s)}{\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})}$
Lemma 4 (Entropy Gradient). For the softmax policy: H(π_θ) = -E_π_θ[_h=0^H-1∇logπ_θ(a_h|s_h)_t=h^H-1 logπ_θ(a_t|s_t)]
Starting from the entropy definition:
$\mathcal{H}(\pi_{\theta})=-\sum_{s_{0},a_{0},\dots,a_{H-1}}\mathbb{P}(s_{0})\prod_{h=0}^{H-1}\pi_{\theta}(a_{h}|s_{h})\sum_{t=0}^{H-1}\log\pi_{\theta}(a_{t}|s_{t})$
Taking the gradient and decomposing:
$\nabla\mathcal{H}(\pi_{\theta})$ $=-\sum_{s_{0},\dots}\mathbb{P}(s_{0})\prod_{h}\pi_{\theta}(a_{h}|s_{h})\nabla\left(\sum_{t}\log\pi_{\theta}(a_{t}|s_{t})\right)$ (26)
$\quad-\sum_{s_{0},\dots}\mathbb{P}(s_{0})\nabla\left(\prod_{h}\pi_{\theta}(a_{h}|s_{h})\right)\sum_{t}\log\pi_{\theta}(a_{t}|s_{t})$ (27)
The first term vanishes because:
$\sum_{a_{0}\dots a_{H-1}}\nabla\left(\prod_{h=0}^{H-1}\pi_{\theta}(a_{h}|s_{h})\right)=\nabla(1)=0$
For the second term, using the product rule and the fact that $\mathbb{E}_{a\sim\pi_{\theta}(s)}[\nabla\log\pi_{\theta}(a|s)]=0$:
$\nabla\mathcal{H}(\pi_{\theta})$ $=-\mathbb{E}_{\pi_{\theta}}\left[\sum_{h=0}^{H-1}\nabla\log\pi_{\theta}(a_{h}|s_{h})\sum_{t=0}^{H-1}\log\pi_{\theta}(a_{t}|s_{t})\right]$ (28)
$=-\mathbb{E}_{\pi_{\theta}}\left[\sum_{h=0}^{H-1}\nabla\log\pi_{\theta}(a_{h}|s_{h})\sum_{t=h}^{H-1}\log\pi_{\theta}(a_{t}|s_{t})\right]$ (29)
The last equality follows from the tower property of expectation, where terms with $t<h$ have zero expectation.

### v2

@@ A.5.4 Performance Bound
Expanding $V_{\lambda,\beta}^{\pi^{*}}$:
$V_{\lambda,\beta}^{\pi^{*}}(s_{0})=V^{\pi^{*}}(s_{0})+\lambda\mathcal{H}(\pi^{*}|s_{0})-\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$
@@
$(I)=V^{\pi^{*}}-V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}\leq V^{\pi^{*}}-V_{\lambda,\beta}^{\pi^{*}}=-\lambda\mathcal{H}(\pi^{*}|s_{0})+\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$ (19)
Step 5: Combining (I) + (III).
$(I)+(III)$ $\leq\lambda\mathcal{H}(\pi_{\theta}|s_{0})-\lambda\mathcal{H}(\pi^{*}|s_{0})+\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$ (20)
$=\lambda(\mathcal{H}(\pi_{\theta}|s_{0})-\mathcal{H}(\pi^{*}|s_{0}))+\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$ (21)
Using entropy bounds: $\mathcal{H}(\pi_{\theta}|s_{0})\leq H\log|\mathcal{A}|$ and $\mathcal{H}(\pi^{*}|s_{0})\geq\log|\mathcal{A}_{H}^{*}(s_{0})|$:
$(I)+(III)\leq\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}+\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$ (22)
Step 6: Final bound.
Combining Steps 2 and 5:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{|\mathcal{D}|^{2}\epsilon^{2}}{2\lambda(1-\alpha\beta_{k})C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}+\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})$ (23)
A.5.5 Zero Stability Cost Condition
See 6.4
Proof.
Part 1: Zero stability cost.
Under the stated condition, $\mathcal{H}(\pi^{*}(\cdot|s))\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$ for all $s\in\mathcal{S}(s_{0})$.
By A.10(2), $\mathcal{P}(\mathcal{H}(\pi^{*}|s);\bar{\mathcal{H}}^{ref})=0$ for all such states.
@@ A.5.5 Zero Stability Cost Condition
$L^{\text{smooth}}_{\bar{\mathcal{H}}^{ref}}(\theta^{*})=\alpha\cdot\mathbb{E}_{s\sim\pi^{*}}[\mathcal{P}(\mathcal{H}(\pi^{*}|s);\bar{\mathcal{H}}^{ref})]=0$
And:
$\Delta^{\text{smooth}}_{\pi^{*}}(s_{0})=\mathbb{E}_{\pi^{*}}[L^{\text{smooth}}|s_{0}]=0$
Part 2: Matching bound.
Substituting $\Delta^{\text{smooth}}_{\pi^{*}}=0$ into 6.3:
$V^{\pi^{*}}-V^{\pi_{\theta}}\leq\frac{|\mathcal{D}|^{2}\epsilon^{2}}{2\lambda(1-\alpha\beta_{k})C}+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}|^{1/H}}$
For small $\alpha$ (e.g., $\alpha=0.1$), the factor $(1-\alpha\beta_{k})^{-1}\approx 1.25$, so the optimization term is only slightly larger than the standard entropy-regularized bound.
Part 3: Additional guarantees.
Even with $\Delta^{\text{smooth}}_{\pi^{*}}=0$, EPO provides stability via A.11: deviations from the corridor are immediately corrected, preventing cascade failure (6.2). ∎
Remark A.18 (When Does the Condition Hold?).
The condition $\mathcal{H}(\pi^{*}|s)\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$ holds when:
  1. 1.
    The task admits multiple near-optimal solutions, requiring stochastic $\pi^{*}$
  2. 2.
    The corridor $\kappa_{l},\kappa_{r}]$ is calibrated to include typical optimal entropy levels
  3. 3.
    Multi-turn tasks where exploration remains valuable throughout the dialogue
In sparse-reward multi-turn settings, deterministic policies often fail to discover good trajectories, so optimal policies naturally maintain moderate entropy.
Remark A.19 (Corridor Width Trade-off).
The corridor bounds $\kappa_{l},\kappa_{r}]$ control a precision-stability trade-off:
Narrow corridor ($\kappa_{r}-\kappa_{l}$ small):
  • Strong entropy enforcement—deviations corrected quickly
  • Higher stability cost if $\pi^{*}$ lies outside corridor
  • Risk: over-constraining may prevent learning optimal behavior
Wide corridor ($\kappa_{r}-\kappa_{l}$ large):
  • Weak entropy enforcement—allows more variation
  • Lower stability cost—most policies satisfy constraint
  • Risk: may not prevent cascade failure if corridor is too permissive
Practical guidance: Set $\kappa_{l},\kappa_{r}]$ to contain the entropy range of good policies while excluding collapse/explosion regimes. Based on experiments across multiple tasks (Appendix 5.3), we find $\kappa_{l}=0.5,\kappa_{r}=1.5$ (allowing entropy to vary by $\pm 50\%$ from reference) provides a robust default.
Sensitivity: Let $w=\kappa_{r}-\kappa_{l}$ denote corridor width. The stability cost scales as:
$\Delta^{\text{smooth}}_{\pi^{*}}\propto\max\left(0,\frac{|\mathcal{H}(\pi^{*})-\bar{\mathcal{H}}^{ref}|}{\bar{\mathcal{H}}^{ref}}-\frac{w}{2}\right)$
Widening the corridor by $\delta$ reduces stability cost by $O(\delta)$ when $\pi^{*}$ is near the boundary.
A.6 Comparison with Standard Entropy Regularization
Corollary A.20 (EPO vs Standard Entropy Regularization).
The standard entropy-regularized bound (shen2025entropy):
$V^{\pi^{*}}-V^{\pi_{\theta}}\leq\frac{|\mathcal{D}|^{2}\epsilon^{2}}{2\lambda C_{\lambda}^{\pi_{\theta}}}+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}|^{1/H}}$
EPO’s bound differs by:
  1. 1.
    A factor of $(1-\alpha\beta_{k})^{-1}\in[1.11,1.25]$ in optimization error (for $\alpha=0.1$)
  2. 2.
    The stability cost $\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}$ (zero under 6.4)
Remark A.21 (Quantifying Multi-Turn Advantage).
For the multi-turn stability advantage (6.2) to outweigh EPO’s per-step overhead, we require:
$\frac{O(T^{2})}{O(T)}=O(T)>\frac{1}{1-\alpha\beta_{k}}$
With $\alpha=0.1$ and $\beta_{k}\in[1,2]$, the right-hand side is at most $1.25$. Thus for $T\geq 2$, the multi-turn stability advantage dominates the per-step overhead. For typical dialogue lengths ($T\geq 5$), the advantage is substantial: $5\times$ to $10\times$ reduction in cumulative entropy deviation.